# Plasmon-assisted photoelectron emission in a model cluster using time-dependent density functional theory and the time–dependent surface–flux method

### v1

@@ IV.2 Scaling of the yield with laser intensity
Figure 7 shows for $\omega_{L}=0.052$, $N_{\text{cyc}}=20$ photoelectron spectra for the three different $a_{0}$ given in the legend. We analyze the ATI peak at energy $E=0.037$, which arises from $\varepsilon_{20}+4\omega_{L}$. Hence we expect an $n=4$ scaling, which is confirmed in Fig. 8 where we find numerically $n=3.88$. We also examine the “plasmon peak” (meaning, “peak due to plasmon-assisted electron emission”) at $E=0.071$, corresponding to $\varepsilon_{19}+2\omega_{B}$. The numerical measurement gives an $n=5.71$-scaling for this peak. One could explain this by observing that it takes three laser photons to excite the $\omega_{B}$-plasmon. The plasmon then acts like a laser itself, and the electron takes one plasmon energy more to escape from the cluster. The two $\omega_{B}$ then correspond to six $\omega_{L}$, which is close to the numerically determined $n=5.71$. However, the electron is emitted when the laser pulse is already over, so the photon energy should not play a role for the emission process. It is rather the strength of the plasmon that should matter for ionization but this is dependent on the laser intensity that excited the plasmon in the first place.
@@
We show another example with the $\omega_{A}$-plasmon involved in Figs. 9 and 10. Again, three photons with $\omega_{L}=0.035$ are required to excite the plasmon. We look at the ATI peak at $\varepsilon_{20}+5\omega_{L}=0.0041$ an find the expected $n=5$-scaling. For the plasmon peak at $\varepsilon_{20}+2\omega_{A}=0.041$ we find numerically $n=6.24\simeq 6$, which also fits with the above interpretation that the excitation of the plasmon scales with $I^{3}$ and the emission process requires another $\omega_{A}=3\omega_{L}$, which compounds to an $I^{6}$-scaling.
@@ V Discussion and outlook
While plasmon-assisted electron emission is clearly present, it remains uncertain whether TDDFT with currently practicable exchange–correlation functionals captures this phenomenon accurately. There are known cases where standard Kohn–Sham TDDFT fails, and more advanced approaches are required Ruggenthaler and Bauer (2009); Dar et al. (2024); Brics and Bauer (2013). Our results indicate that plasmon-assisted electron emission is likely overemphasized in the present simulations: the associated oscillations should decay more rapidly after the laser pulse. This behavior can be attributed to several limitations of our model. First, Landau damping is strongly suppressed in one dimension. In higher dimensions, the larger density of states would enhance dephasing, leading to a shorter plasmon lifetime. Second, real dissipation channels, such as coupling to phonons or the environment, are absent from our description. Third, the exact exchange–correlation potential, if it were available, would likely accelerate the decay of the plasma oscillations. Plasmon-assisted electron emission may therefore serve as a stringent test case for the development of improved exchange–correlation functionals. Furthermore, it remains unclear whether the scaling of the plasmonic peaks in the photoelectron spectra is captured quantitatively by our TDDFT simulations. There are other examples in strong-field physics where TDDFT predicts qualitatively correct but quantitatively inaccurate features, such as the second plateau in high-harmonic generation Brics et al. (2016). Future work will thus focus on models amenable to higher-level many-body techniques, allowing for systematic benchmarking of TDDFT in this challenging regime.
@@ Acknowledgements.
We acknowledge funding from the Deutsche Forschungsgemeinschaft via SFB 1477 ’Light–Matter Interactions at Interfaces’ (project no. 441234705).

### v2

@@ IV.2 Scaling of the yield with laser intensity
Figure 7 shows for $\omega_{L}=0.052=\omega_{B}/3$, $N_{\text{cyc}}=20$ photoelectron spectra for the three different $a_{0}$ given in the legend. We analyze the ATI peak at energy $E=0.037$, which arises from $\varepsilon_{20}+4\omega_{L}$. Hence we expect an $n=4$ scaling, which is confirmed in Fig. 8 where we find numerically $n=3.88$. We also examine the “plasmon peak” (meaning, “peak due to plasmon-assisted electron emission”) at $E=0.071$, corresponding to $\varepsilon_{19}+2\omega_{B}$. The numerical measurement gives an $n=5.71$-scaling for this peak. One could explain this by observing that it takes three laser photons to excite the $\omega_{B}$-plasmon. The plasmon then acts like a laser itself, and the electron takes one plasmon energy more to escape from the cluster. The two $\omega_{B}$ then correspond to six $\omega_{L}$, which is close to the numerically determined $n=5.71$.
@@
We show another example with the $\omega_{A}$-plasmon involved in Figs. 9 and 10. Again, three photons with $\omega_{L}=0.035$ are required to excite the plasmon. We look at the ATI peak at $\varepsilon_{20}+5\omega_{L}=0.0041$ and find the expected $n=5$-scaling. For the plasmon peak at $\varepsilon_{20}+2\omega_{A}=0.041$ we find numerically $n=6.24\simeq 6$, which also fits with the above interpretation that the excitation of the plasmon scales with $I^{3}$ and the emission process requires another $\omega_{A}=3\omega_{L}$, which compounds to an $I^{6}$-scaling.
@@ V Discussion and outlook
While plasmon-assisted electron emission exists, it remains uncertain whether TDDFT with currently practicable exchange-correlation functionals captures this phenomenon accurately. There are known cases where standard Kohn-Sham TDDFT fails, and more advanced approaches are required Ruggenthaler and Bauer (2009); Dar et al. (2024); Brics and Bauer (2013). Our results indicate that plasmon-assisted electron emission is likely overemphasized in the present simulations: the associated oscillations should decay more rapidly after the laser pulse. This behavior can be attributed to several limitations of our model. First, Landau damping is suppressed in one dimension. In higher dimensions, the larger density of states would enhance dephasing, leading to a shorter plasmon lifetime. Second, real dissipation channels, such as coupling to phonons or the environment, are absent from our description. Third, the exact exchange-correlation potential, if it were available, would likely accelerate the decay of the plasma oscillations. Plasmon-assisted electron emission may therefore serve as a stringent test case for the development of improved exchange-correlation functionals. Furthermore, it remains unclear whether the scaling of the plasmonic peaks in the photoelectron spectra is captured quantitatively by our TDDFT simulations. There are other examples in strong-field physics where TDDFT predicts qualitatively correct but quantitatively inaccurate features, such as the second plateau in high-harmonic generation Brics et al. (2016). Future work will thus focus on models amenable to higher-level many-body techniques, allowing for systematic benchmarking of TDDFT in this challenging regime.
@@ Acknowledgements.
We acknowledge funding from the Deutsche Forschungsgemeinschaft via SFB 1477 ’Light-Matter Interactions at Interfaces’ (project no. 441234705).